Suppose that the mean score of a class of 35 students was 76. The 17 male students in the class had a mean score of 62. What was the mean score of the 18 female students?
Answer by stanbon(75887)

Suppose that the mean score of a class of 35 students was 76. The 17 male students in the class had a mean score of 62. What was the mean score of the 18 female students?
------------------
Total class points = 35*76
Total male points = 17*62
Total female points = (35*76-17*62) = 1606
-------------------
Mean female score = 1606/18 = 89.2222...
Cheers,
Stan H.